# supplemental

### v2

@@ A.2.1 Mask as Input or Output
”Marbel
mosaic”
”Ice”
”Leather”
Figure 12: Evaluation of our model’s robustness to varying levels of distortion. The first column shows textures transformed with homography and thin plate spline transformations. The following columns present the material maps and the rendering images. The results demonstrate that our model effectively rectifies textures with various patterns and different types of distortion, maintaining high-quality outputs.
”Wood”
@@
”Leather”
”Marble”
Figure 13: Evaluation of the robustness of our model to lighting and shadow interference. We test scenarios where the input photographs contain point light sources, shadows, or environmental reflections. The generated material maps and rendered images demonstrate the ability of our model to handle these interferences, preserving material quality and accurately representing the input photos. The leftmost side of each row is labeled with the text conditioning input used.
Appendix B Additional Real Examples
In Fig. 14, we provide the material maps of the five examples used for texture synthesis in Fig. 5 of the main paper. In Fig. 17, Fig. 18, Fig. 19, Fig. 20, Fig. 21 and Fig. 22, we show additional examples of material extraction using our model, along with the uncropped original images from our real photographs evaluation dataset (Sec. 3.2). These examples include various indoor and outdoor materials captured under complex real-world lighting conditions. Our model generalizes well to real photos, producing renderings that are visually similar to the photographs and providing accurate masks, demonstrating our model’s generalization capabilities.
@@ Appendix B Additional Real Examples
Figure 14: The material maps of the five samples used for texture rectification comparison in Fig. 5 of the main paper.
@@
Figure 15: More results of comparisons with Material palette [28] on synthetic dataset (Sec. 4.1) for material extraction. Each row contains two sets of comparisons. In each set, the first column shows the ground truth material maps from PolyHaven, with the rendered scene below. The yellow square area indicates the crop used as the input for both models. The second and third columns show the material maps extracted by our model and Material Palette, along with the re-rendered images. We can see that our approach better matches the Ground-Truth appearance.
@@
Figure 16: More results of comparisons with Material palette [28] on synthetic dataset (Sec. 4.1) for material extraction. Each row contains two sets of comparisons. In each set, the first column shows the ground truth material maps from PolyHaven, with the rendered scene below. The yellow square area indicates the crop used as the input for both models. The second and third columns show the material maps extracted by our model and Material Palette, along with the re-rendered images. We can see that our approach better matches the Ground-Truth appearance.
@@
Figure 17: More results of our method on material extraction for real photographs. The first column shows real photographs captured by smartphones, before cropping (as described in Sec. 3.2). The box indicates the cropped area, which is the image actually used as input for the model in the second column. Third to the ninth columns show the generated material maps and rendering under two environment maps. The last column shows the mask of the dominant material location automatically predicted by our model. We conduct tests on diverse material types in both indoor and outdoor scenes, demonstrating the generalization capability of our model. The leftmost side of each row is labeled with the text conditioning input used.
@@
Figure 18: More results of our method on material extraction for real photographs. The first column shows real photographs captured by smartphones, before cropping (as described in Sec. 3.2). The box indicates the cropped area, which is the image actually used as input for the model in the second column. Third to the ninth columns show the generated material maps and rendering under two environment maps. The last column shows the mask of the dominant material location automatically predicted by our model. We conduct tests on diverse material types in both indoor and outdoor scenes, demonstrating the generalization capability of our model. The leftmost side of each row is labeled with the text conditioning input used.
@@
Figure 19: More results of our method on material extraction for real photographs. The first column shows real photographs captured by smartphones, before cropping (as described in Sec. 3.2). The box indicates the cropped area, which is the image actually used as input for the model in the second column. Third to the ninth columns show the generated material maps and rendering under two environment maps. The last column shows the mask of the dominant material location automatically predicted by our model. We conduct tests on diverse material types in both indoor and outdoor scenes, demonstrating the generalization capability of our model. The leftmost side of each row is labeled with the text conditioning input used.
@@
Figure 20: More results of our method on material extraction for real photographs. The first column shows real photographs captured by smartphones, before cropping (as described in Sec. 3.2). The box indicates the cropped area, which is the image actually used as input for the model in the second column. Third to the ninth columns show the generated material maps and rendering under two environment maps. The last column shows the mask of the dominant material location automatically predicted by our model. We conduct tests on diverse material types in both indoor and outdoor scenes, demonstrating the generalization capability of our model. The leftmost side of each row is labeled with the text conditioning input used.
@@
Figure 21: More results of our method on material extraction for real photographs. The first column shows real photographs captured by smartphones, before cropping (as described in Sec. 3.2). The box indicates the cropped area, which is the image actually used as input for the model in the second column. Third to the ninth columns show the generated material maps and rendering under two environment maps. The last column shows the mask of the dominant material location automatically predicted by our model. We conduct tests on diverse material types in both indoor and outdoor scenes, demonstrating the generalization capability of our model. The leftmost side of each row is labeled with the text conditioning input used.
@@
Figure 22: More results of our method on material extraction for real photographs. The first column shows real photographs captured by smartphones, before cropping (as described in Sec. 3.2). The box indicates the cropped area, which is the image actually used as input for the model in the second column. Third to the ninth columns show the generated material maps and rendering under two environment maps. The last column shows the mask of the dominant material location automatically predicted by our model. We conduct tests on diverse material types in both indoor and outdoor scenes, demonstrating the generalization capability of our model. The leftmost side of each row is labeled with the text conditioning input used.

### v3

@@ 4.4.4. Input scale
"Marble tiles"
"Marble tiles"
Figure 10. Evaluation of our model’s adaptability to different input texture scales on real photographs. We can see that our results are generated with a scale matching that of the input.
4.4.5. Evaluations on the Robustness.
To examine the robustness of our model to strong, real-world, distortions, we generate a synthetic test set that use textures from the texture datasets TexSD (Lopes et al., 2024) and follow the texture processing steps outlined by Hao et al. (2023). We apply homography transformations (Hartley and Zisserman, 2003) and thin plate spline transformations (Bookstein, 1989) to the textures. Our results in Fig. 11 show that the model is robust to severe distortions, stretching, and the blurring effects introduced by these transformations. More examples of real photos with distortion or surface geometry diversity can be found in supplemental materials.
"Gingham"
@@
"Marble
mosaic"
"Ice"
"Leather"
Figure 11. Evaluation of our model’s robustness to varying levels of distortion. The first column shows textures transformed with homography and thin plate spline transformations. The following columns present the material maps and the rendering images. The results demonstrate that our model effectively rectifies textures with various patterns and different types of distortion, maintaining high-quality outputs.
"Wood"
@@
"Leather"
"Marble"
Figure 12. Evaluation of the robustness of our model to lighting and shadow interference. We test scenarios where the input photographs contain point light sources, shadows, or environmental reflections. The generated material maps and rendered images demonstrate the ability of our model to handle these interferences, preserving material quality and accurately representing the input photos. The leftmost side of each row is labeled with the text conditioning input used.
To further demonstrate the generalization ability of our model beyond common indoor and outdoor scenes, we test several images sourced from Pixabay [^2] https://pixabay.com/ (royalty-free). The results are shown in Fig. 13. Notably, for challenging appearances such as animal skin and fur, and plant surfaces, our model can generate visually plausible results.
"Snake skin"
(a) Input
(b) Albedo
(c) Normal
(d) Roughness
(e) Height
(f) Metallic
(g) Render
"Frog skin"
"Tiger fur"
"Cat fur"
"Strawberry"
"Leaves"
"Moon
surface"
"Manhole
cover"
Figure 13. Evaluation of the generalization ability of our model to complex patterns. We test on several royalty-free images from Pixabay, including animal and plant appearances as well as abstract textures like the moon surface. The results demonstrate that our model generalizes well beyond typical indoor and outdoor scenes, producing visually plausible outputs for a diverse range of complex surface patterns.
4.4.6. Tileable Generation.
Although our model is not explicitly trained to produce tileable outputs, we can easily incorporate noise rolling, a test-time technique proposed in ControlMat (Vecchio et al., 2024a), to generate seamless material maps without any re-training. We present several successful examples of applying this technique in Fig. 14.
(a) Input
(b) 512$\times$512
(c) 1024$\times$1024
(d) Render
"checkered
fabric"
"Wood"
"Terracotta
brick wall"
"Stone path"
Figure 14. Evaluation of tileable material generation. During inference, we can employ the noise rolling strategy (Vecchio et al., 2024a) to enhance the seamlessness of the generated results. The first column shows the input images, followed by the 512$\times$512 material maps. The third column displays 1024$\times$1024 tiling results obtained by stitching the 512-resolution outputs, and the final column presents the rendering results using the 1024-resolution maps. Maps are ordered as albedo, normal, roughness, and height, from top to bottom and left to right. Metallic maps are omitted as they are black in these cases.
5. Limitations
Despite strong generation capacity, our model may still encounter challenging inputs, as shown in Fig. 15. In the first row we show an example where our model confuses shading and albedo variation. Our model may also have difficulty handling materials with cutouts or holes, since it does not produce opacity maps as outputs. Also, preserving semantically meaningful patterns, such as text, is a remaining challenge in our approach. Furthermore, our current model is not capable of handling transparency and refraction effects. In the third row of Fig. 11, we present an example of generating "ice", where the model tries to approximates the appearance using the albedo map. The fourth and fifth rows of Fig. 15 illustrate more challenging cases, in which the model fails to reproduce the visual characteristics of transparent and translucent materials due to the lack of relevant training data.
To improve the model’s capacity for handling complex appearance effects such as transparency, translucency, and detailed BRDF components, one promising direction is to incorporate more advanced material maps (e.g., specular, coat, and subsurface scattering) into the training data, illustrated in Fig. 16. As outlined in Sec. 3.3, our architecture is flexible and can be extended to generate additional output channels corresponding to these effects. We leave this extension for future work, once sufficiently large-scale datasets with such complex annotations become available.
"Concrete
pavement"
(a) Input
(b) Albedo
(c) Normal
(d) Roughness
(e) Height
(f) Metallic
(g) Render
"Perforated
metal"
"Towel"
"Glass"
"Sheer
curtain "
Figure 15. Limitations. We show limitations of our model, such as complex lighting and shadows in the first row, materials with perforations in the second row, structurally significant elements like text in the third row, and effects involving transparency and translucency in the last two rows.
Figure 16. Future work. These images are obtained using the method described in Sec. 3.2, with the addition of a translucency map during rendering. We consider such complex scenarios as future work.
6. Conclusion
We present a generative model for high-quality material synthesis from text prompts and/or crops of natural images by finetuning a pretrained text-to-video generative model, which provides strong prior knowledge. The flexible video DiT architecture lets us adjust the model for multi-channel material generation. We show extensive evaluation on both synthetic and real examples and conduct systematic ablation studies and test on robustness. We believe that our re-purposing of a video model for multi-channel generation opens an interesting avenue for other domain which require the generation of additional channels, such as intrinsic decomposition (Vecchio et al., 2024a).
(a) Input
(b) Mask (Input)
(c) [Hao et al. 2023]
(d) Our Mask (Output)
(e) Ours (Render 1)
(f) Ours (Render 2)
(g) Albedo
(h) Normal
(i) Roughness
(j) Height
(k) Metallic
Figure 17. Comparison with Hao et al. (2023) on texture rectification for real photos. The first column shows the input photos. The second and third columns are the required input masks and output textures of Hao et al. The fourth column shows masks generated by our model, followed by two renderings (Render 1 & Render 2) of our generated material maps (in the last five columns) under different environment maps. Despite not requiring an input mask, our method better rectifies perspective and distortions. Further, as we support material extraction, our result does not contain shading from the input image.
Appendix A More synthetic examples
We show qualitative and quantitative comparisons with Material Palette (Lopes et al., 2024) on material extraction in Sec. 4.2 of the main paper. We provide more visual comparisons in Fig. 18 and Fig. 19. Since Material Palette generates only three material maps (albedo, normal, and roughness), we present results for these channels, along with the re-rendered images.
Appendix B More Real Examples
In Fig. 20, Fig. 21, Fig. 22, Fig. 23, Fig. 24 and Fig. 25, we show additional examples of material extraction using our model, along with the uncropped original images from our real photographs evaluation dataset (Sec. 3.2). These examples include various indoor and outdoor materials captured under complex real-world lighting conditions. Our model generalizes well to real photos, producing renderings that are visually similar to the photographs and providing accurate masks, demonstrating our model’s generalization capabilities.
@@ Appendix B More Real Examples
Figure 18. More results of comparisons with Material palette (Lopes et al., 2024) on synthetic dataset (Sec. 4.1) for material extraction. Each row contains two sets of comparisons. In each set, the first column shows the ground truth material maps from PolyHaven, with the rendered scene below. The yellow square area indicates the crop used as the input for both models. The second and third columns show the material maps extracted by our model and Material Palette, along with the re-rendered images. We can see that our approach better matches the Ground-Truth appearance.
@@
Figure 19. More results of comparisons with Material palette (Lopes et al., 2024) on synthetic dataset (Sec. 4.1) for material extraction. Each row contains two sets of comparisons. In each set, the first column shows the ground truth material maps from PolyHaven, with the rendered scene below. The yellow square area indicates the crop used as the input for both models. The second and third columns show the material maps extracted by our model and Material Palette, along with the re-rendered images. We can see that our approach better matches the Ground-Truth appearance.
@@
Figure 20. More results of our method on material extraction for real photographs. The first column shows real photographs captured by smartphones, before cropping (as described in Sec. 3.2). The box indicates the cropped area, which is the image actually used as input for the model in the second column. Third to the ninth columns show the generated material maps and rendering under two environment maps. The last column shows the mask of the dominant material location automatically predicted by our model. We conduct tests on diverse material types in both indoor and outdoor scenes, demonstrating the generalization capability of our model. The leftmost side of each row is labeled with the text conditioning input used.
@@
Figure 21. More results of our method on material extraction for real photographs. The first column shows real photographs captured by smartphones, before cropping (as described in Sec. 3.2). The box indicates the cropped area, which is the image actually used as input for the model in the second column. Third to the ninth columns show the generated material maps and rendering under two environment maps. The last column shows the mask of the dominant material location automatically predicted by our model. We conduct tests on diverse material types in both indoor and outdoor scenes, demonstrating the generalization capability of our model. The leftmost side of each row is labeled with the text conditioning input used.
@@
Figure 22. More results of our method on material extraction for real photographs. The first column shows real photographs captured by smartphones, before cropping (as described in Sec. 3.2). The box indicates the cropped area, which is the image actually used as input for the model in the second column. Third to the ninth columns show the generated material maps and rendering under two environment maps. The last column shows the mask of the dominant material location automatically predicted by our model. We conduct tests on diverse material types in both indoor and outdoor scenes, demonstrating the generalization capability of our model. The leftmost side of each row is labeled with the text conditioning input used.
@@
Figure 23. More results of our method on material extraction for real photographs. The first column shows real photographs captured by smartphones, before cropping (as described in Sec. 3.2). The box indicates the cropped area, which is the image actually used as input for the model in the second column. Third to the ninth columns show the generated material maps and rendering under two environment maps. The last column shows the mask of the dominant material location automatically predicted by our model. We conduct tests on diverse material types in both indoor and outdoor scenes, demonstrating the generalization capability of our model. The leftmost side of each row is labeled with the text conditioning input used.
@@
Figure 24. More results of our method on material extraction for real photographs. The first column shows real photographs captured by smartphones, before cropping (as described in Sec. 3.2). The box indicates the cropped area, which is the image actually used as input for the model in the second column. Third to the ninth columns show the generated material maps and rendering under two environment maps. The last column shows the mask of the dominant material location automatically predicted by our model. We conduct tests on diverse material types in both indoor and outdoor scenes, demonstrating the generalization capability of our model. The leftmost side of each row is labeled with the text conditioning input used.
@@
Figure 25. More results of our method on material extraction for real photographs. The first column shows real photographs captured by smartphones, before cropping (as described in Sec. 3.2). The box indicates the cropped area, which is the image actually used as input for the model in the second column. Third to the ninth columns show the generated material maps and rendering under two environment maps. The last column shows the mask of the dominant material location automatically predicted by our model. We conduct tests on diverse material types in both indoor and outdoor scenes, demonstrating the generalization capability of our model. The leftmost side of each row is labeled with the text conditioning input used.